OpenStudy (anonymous):

Explain this: S/(S²+cos²) = sec²Θ / (S²sec²Θ + 1) = sec²Θ / [(S²+1)+S²tan²Θ]

OpenStudy (akashdeepdeb):

\[\frac{S}{(S^2 + cos^2 \theta)}\] = \[\frac{S}{S^2 + \frac{1}{sec^2 \theta}}\] Because \(cos^2 \theta = \frac{1}{sec^2 \theta}\) Then \[\frac{S*sec^2 \theta}{sec^2 \theta *S^2 + 1}\] = \[\frac{S*sec^2 \theta}{S^2 *(1 + tan^2 \theta) + 1}\] Because \(1 + tan^2 \theta = sec^2 \theta\) And then = \[\frac{S*sec^2 \theta}{(S^2+1) + S^2*tan^2 \theta}\] Understood? :)

OpenStudy (anonymous):

Mother of explanation..... Thank you !!!

OpenStudy (akashdeepdeb):

:)